OpenStudy (anonymous):

sin2x cosx + cos 2x sin x= -1

OpenStudy (zzr0ck3r):

that is sin(2x+x) = 1

OpenStudy (zzr0ck3r):

can you solve that?

OpenStudy (zzr0ck3r):

woops -1 sorry so sin(2x+x) = sin(3x) = -1 so 3x = -pi//2

OpenStudy (zzr0ck3r):

I think someone else is about to same the same thing but write it cooler:)

OpenStudy (jhannybean):

LOL -_-

OpenStudy (zzr0ck3r):

ha

OpenStudy (jhannybean):

Tbh i'm using identities and i don't know if it's making it more complicated, lol..

OpenStudy (zzr0ck3r):

I used sin(a+b) = sin(a)cos(b) + cos(a)sin(b)

OpenStudy (jhannybean):

\[\sin(2x)\cos(x)+\cos(2x)\sin(x)=-1\]\[2\sin(x)\cos(x)*\cos(x)+(1-2\sin^2(x))*\sin (x)=-1\]\[2\sin(x)\cos^2(x)-\sin(x)-2\sin^3(x)=-1\]\[\sin(x)(2\cos^2(x)-1-2\sin^2(x))=-1\]\[\sin(x)(\cos(2x)-2\sin^2(x))=-1\] Yeah no....

OpenStudy (zzr0ck3r):

sin(3x) = sin(2x+x) = sin(2x)cos(x) + cos(2x)sin(x)

OpenStudy (jhannybean):

Oh is that right...lmao ok.